Normalize the function Ψ = e^-r

1
Expert's answer
2022-03-29T16:46:24-0400

"\int_0^{\infin}a^2e^{-2r}dr=1,"

"\frac{a^2}{-2}e^{-2r}|_0^{\infin}=1,"

"\frac{a^2}2(1-0)=1,"

"a=\sqrt 2,"

"\psi=\sqrt2e^{-r}."